An electromagnetic wave strikes a 1.48-cm2 section of wall perpendicularly. The rms value of the wave's magnetic field is determined to be 6.27 × 10-4 T. How long does it take for the wave to deliver 2190 J of energy to the wall?

Will Rate
Answer #1

The max value of the wave's magnetic field is:

The Poynting vector is given by :
S = cB² / µ₀
S = ((3.0 * 10^8) * (8.87 * 10^-4)^2) / (4π * 10^-7)
S = 9.385 * 10^7 W/m²

The power delivered to the wall is:

P = SA = (9.385 * 10^7 W/m²) x (1.48 * 10^-4 m²)
P = 13889.8 W

The time needed is:

t = E / P = 2190 J / 13889.8 W
t = 0.158 s